OpenStudy (anonymous):

How do you find an equation of the line containing the given pair of points (-7,-3) and (-2,-1)

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\] to find your slope, then after you do that pick one point you're given and plug it into y=mx+b along with your slope to find the y-intercept, then write your equation

OpenStudy (anonymous):

The difference between the two points is the slope. \[\frac{ -1 - (-3) }{ -2 - (-7) }\] \[\frac{ -1+3 }{ -2+7 }\] \[\frac{ 2 }{ 5 }\] With the slope, you can use point-slope form for the equation of a line. \[y-y_{1}=m(x-x_{1})\] \[y+1=\frac{ 2 }{ 5 } (x+2)\]